OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 9x + 10 = 0? (3 points) No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

b^-4ac= 81-4x1x10 =1 hence two different rational solutions

OpenStudy (anonymous):

calculate b^2-4ac if it is positive then you have two distinct roots if it is zero then you have two identical roots if it is negative then you have two imaginay roots

OpenStudy (anonymous):

so its the third one then? @irshadno1

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

ok, awesome!!

OpenStudy (anonymous):

b^-4ac= 81-4x1x10 = 81- 40 = 41 it is positive so you have two distinct roots

OpenStudy (anonymous):

ok, what does that mean?

OpenStudy (anonymous):

yes u r right ....